DIRECT VARIATION
Find an equation of variation where y varies directly as x and y=2 when x=10 find y when x = -25.
y=____
Thanks you very much

Answer by Nate(3500)

Find an equation of variation where y varies directly as x and y=2 when x=10 find y when x = -25.
(y)/(x)
(2)/(10) = (y)/(-25)
-5 = y